OpenStudy (anonymous):

12 sqrt (-3)^12 Simplify your answer.

OpenStudy (anonymous):

12 sqrt (-3)^12 12 sqrt (531,441) 12*729 8748.

OpenStudy (anonymous):

12(-3)^6 =8748

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt[12]{(-3)^{12}}=|-3|=3\] So \[\Large \sqrt[12]{(-3)^{12}}=3\]

OpenStudy (anonymous):

12(-3)^6 =8748

OpenStudy (anonymous):

Raheen and Royal, what do you think about Jim's response?

jimthompson5910 (jim_thompson5910):

I'm assuming that 12 sqrt refers to the 12th root

OpenStudy (anonymous):

Yes, it does. You wrote the problem out correctly. 3 other people had the same answer, so I'm wondering if they misunderstood the way i wrote the proble, not that I'm questioning your answer. (*.*)

jimthompson5910 (jim_thompson5910):

ok, then the answer I have is correct they're thinking it's 12 times the square root of (-3)^12

OpenStudy (anonymous):

ok, thanks for the clarity!

OpenStudy (anonymous):

Yes, I thought it was 12 times the square root of (-3)^12. Sorry. Jim, how do you know if it refers to 12th root? Just curious, I am still learning! :)

jimthompson5910 (jim_thompson5910):

i've just seen problems like this for a long time, so practice I guess

OpenStudy (anonymous):

So Jim...DId I type it incorrectly, and if so how should I have typed it?

jimthompson5910 (jim_thompson5910):

type either root(12,(-3)^12) or 12th root of (-3)^12 If in doubt or you forget, then just verbally describe it (eg: "There's a square root with a little 12 on top of the square root up to the left. Under the square root is (-3)^12"). Sure that may seem a bit wordy, but it's probably the most natural way to do it.

jimthompson5910 (jim_thompson5910):

or you can draw it (probably a better way)